A frog spots a Lily pad in the water 2.50 m from the shore. He jumps off a tree branch 44 em above the shore at an angle of 35° above the horizontal with an initial velocity of 2.4 m/s. If the Lily pad is.12 em in diameter does he make his target or does he land in the water? No points for guessing. You must show your work 4.
Answer #1

in vertical,

v0y = 2.4 sin35 = 1.38 m/s

ay = - 9.8 m/s^2

displacement, y = 0 - 0.44 = - 0.44 m

Applying y = v0y t + ay t^2 /2

4.9 t^2 - 1.38t - 0.44 = 0

t = 0.47 sec

in horizontal,

d = (v0x)(t)

d = (2.4 cos35) (0.47)

d = 0.92 m


so it will not reach to the lily pad.

He land in water. ...Ans